# Adversary Resilient Learned Bloom Filters

### v3

@@ 4 Discussion
In this section, we discuss the only two known attacks on the Learned Bloom Filter, introduced by Reviriego et al [RAHDS21]. We refer to Attack 1 from their work as the Blackbox Mutation Attack, and Attack 2 from their work as the Whitebox Mutation Attack respectively. We discuss the Blackbox Mutation Attack in Section 4.1, and the Whitebox Mutation Attack in Section 4.2. Both sections include details on how our Secure Learned Bloom Filter construction mitigates the attack.
@@ 4.1 Black-box Mutation Attack
The black-box adversarial model defined by Reviriego et al [RAHDS21] is slightly weaker but very similar to the adversarial model we define in Challenge 3 of Section 2.2. Both our adversarial model and Reviriego et al’s black-box adversary model allows the adversary access to query the Learned Bloom Filter. One major difference is that our model allows the adversary to choose the initial set $S$ that is represented by the Learned Bloom Filter, whereas Reviriego et al’s model does not. In their attack, Reviriego et al first test elements until a positive (whether a false positive or true positive) is found. They then mutate the positive by changing a small fraction of the bits in the input in order to generate more false positives. The attack targets the Bloom Filter Learning Model (recall Figure 2) by making it generate false positives without the input reaching the Backup Bloom Filter. This attack is mitigated by the Downtown Bodega Filter because the Downtown Bodega Filter passes queries through a Secure Classical Backup Bloom Filter even for the case where the Bloom Filter Learning Model returns true i.e when $l(x)>\tau$.
@@ 4.2 White-box Mutation Attack
The white-box adversarial model defined by Reviriego et al [RAHDS21] is similar to the adversarial model we define in Challenge 4 in Section 2.2. With knowledge of the state of the Bloom Filter Learning Model, the adversary can generate mutations in a more sophisticated way. Reviriego et al provide the example of a malicious URL dataset where an adversary may begin with a non-malicious URL and make changes such as removing the “s” in “https” or removing the “www” to generate false positives. Since we have shown the Downtown Bodega Filter to be $(n,t,\epsilon)$-adversarial reveal resilient in the steady setting, such mutations will not provide the adversary any advantage over the construction.
@@ 5 Hybrid Adversarial Model
In this section, we first define a hybrid model, where part of the queries are chosen by an adversary, while the rest are non-adversarial (“regular”) queries. Next we analyze the performance of the Downtown Bodega Filter and the Secure Classical Bloom Filter respectively in our hybrid model. We provide results for the conditions in which the Downtown Bodega Filter outperforms the Secure Classical Bloom Filter in the hybrid model. We then discuss the trade-offs of our approach and provide a realistic example where the Downtown Bodega Filter displays better performance as compared to the Secure Classical Bloom Filter under many settings.
@@ 5.1 Hybrid Model
Let $A=(A_{C},A_{Q})$ be a probabilistic polynomial time (PPT) adversary as defined in Section 2. Consider a set of $N$ queries sent to a (Classical or Learned) Bloom Filter. Our adversary, $A$, is allowed to choose exactly $\alpha N$ of those queries, where $\alpha\in[0,1]$. As an example, the $N$ queries may be part of a streaming workload under any of the streaming models described by Muthukrishnan [Mut05].
@@ 5.2 Downtown Bodega Filter
Let $DB_{r}=(\tilde{C}_{r},\tilde{Q}_{s})$ be a Downtown Bodega Filter. Let $l_{r}$ be the corresponding Bloom Filter Learning Model and $\tau$ its threshold. Let $M_{r}^{S^{\prime\prime}}$ be the internal representation of Backup Classical Bloom Filter $A$ and $M_{r}^{S^{\prime}}$ be the internal representation of Backup Classical Bloom Filter $B$. For an adversarial query to generate a false positive in the Downtown Bodega Filter, one of the following must hold true for a given query (see Figure 5):
@@
    The query generates a false positive in the Bloom Filter Learning Model and a false positive in Backup Bloom Filter $A$.
@@
    The query generates a true negative in the Bloom Filter Learning Model and a false positive in Backup Bloom Filter $B$
@@
Let $\text{FPR}(S,m)$ be the expected false positive probability of a Classical Bloom Filter that encodes the set $S$ with memory budget $m$. Let $\text{FPR}_{L}(S,\mathcal{D},m)$ be the expected false positive probability of a Bloom Filter Learning Model $L$ that encodes the set $S$ using the training dataset $\mathcal{D}$ with memory budget $m$. Similarly, let $\text{TNR}_{L}(S,\mathcal{D},m)$ be the expected true negative probability of a Bloom Filter Learning Model $L$ that encodes the set $S$ using the training dataset $\mathcal{D}$ with memory budget $m$. We make the assumption that the correctness probability of the Bloom Filter Learning Model is independent of the correctness probability of the Backup Classical Bloom Filters. In particular, we assume that for any $m,m^{\prime}$, $\text{FPR}_{L}(m)\cap\text{FPR}(m^{\prime})=\text{FPR}_{L}(m)\text{FPR}(m^{\prime})$ and $\text{TNR}_{L}(m)\cap\text{FPR}(m^{\prime})=\text{TNR}_{L}(m)\text{FPR}(m^{\prime})$.
Consider a system where the total memory budget is $M$. Let the memory allocation of a Downtown Bodega Filter from memory budget $M$ be assigned as follows. Let $m_{L}$ be the number of bits of memory assigned to a Bloom Filter Learning Model $L$. Let $m_{A}$ be the number of bits of memory assigned to Backup Bloom Filter $A$. Let $m_{B}$ be the number of bits of memory assigned to Backup Bloom Filter $B$. Let $\lambda$ be the number of bits assigned to the key of the pseudo-random permutations used prior to Backup Bloom Filter $A$ and Backup Bloom Filter $B$. Note that to stay within the memory budget it must hold that $M\geq m_{L}+m_{A}+m_{B}+2\lambda$.
Theorem 5.1.
@@
Proof 5.2.
From the definition of a Downtown Bodega Filter (Definition 3.4, in particular look at the formulation for the query algorithm $\tilde{Q}_{s}$. Figure 5 is also helpful here) it follows that a false positive in the overall construction must either be a false positive in Backup Classical Bloom Filter $A$ or a false positive in Backup Classical Bloom Filter $B$. If the query is a false positive in Backup Classical Bloom Filter $A$, it must also be a false positive in the Bloom Filter Learning Model. Alternatively, if the query is a false positive in Backup Classical Bloom Filter $B$, it must be a true negative in the Bloom Filter Learning Model. The result follows.
@@
Case 1: The query is not an adversary generated query. Therefore as established by Theorem 5.1, the False Positive Probability for the query is $\text{FPR}_{DB}(S,\mathcal{D},M)$. There are $(1-\alpha_{P}-\alpha_{N})N$ such queries.
Case 2: The query is an adversary generated query such that it generates a false positive in the Bloom Filter Learning Model. Since the Bloom Filter Learning Model generating a false positive and the Bloom Filter Learning Model generating a true negative are mutually exclusive events, the False Positive Probability for the query is the False Positive Probability of Backup Classical Bloom Filter $A$ i.e $\text{FPR}(S^{\prime\prime},m_{A})$. There are $\alpha_{P}N$ such queries.
Case 3: The query is an adversary generated query such that it generates a true negative in the Bloom Filter Learning Model. Following logic similar to case 2, we can derive the False Positive Probability of the query to be $\text{FPR}(S^{\prime},m_{B})$. There are $\alpha_{N}N$ such queries.
@@ 5.3 Secure Classical Bloom Filter
An alternative construction is to simply use a well-tuned $(n,t,\epsilon)$-adversarial reveal resilient Classical Bloom Filter. We will refer to this construction in this section as the Secure Classical Bloom Filter without causing confusion. The expected false positive probability of the Secure Classical Bloom Filter encoding a set $S$ is, by definition, $\text{FPR}(S,m_{L}+m_{A}+m_{B}+\lambda)$. The extra $\lambda$ is due to the fact that the Secure Classical Bloom Filter requires one less pseudo-random permutation than the Downtown Bodega Filter.
@@
Theorem 5.5.
For a given memory budget $M=m_{L}+m_{A}+m_{B}+2\lambda$, any set $S$ and corresponding training dataset $\mathcal{D}$, in the hybrid adversarial setting with given $\alpha=\alpha_{P}+\alpha_{N}$, the expected false positive probability of the Downtown Bodega Filter is lower than the expected false positive probability of the Secure Classical Bloom Filter if the following holds true:
@@
5.4 Realistic Example
Mitzenmacher Broder [MB04] show that the false positive probability for a Classical Bloom Filter with $m$ bits encoding a set $S$, using $k(S,m)$ hash functions is
$\textrm{FPR}(S,m)=(1-e^{-k(S,m)\cdot|S|/m})^{k(S,m)}$
In our analysis, we use the value of the number of hash functions $k(S,m)$ is always optimally chosen to be $k(S,m)=\ln 2\cdot(m/|S|)$ (this optimal value is also derived by Mitzenmacher Broder [MB04]). We further model the false positive rate of a Bloom Filter Learning Model as being the same as the false positive rate of a Classical Bloom Filter encoding set $S$ but with a better false positive probability for the same memory budget. This is consistent with the assumptions made by prior work including Kraska et al [KBC+18] and Mitzenmacher [Mit18].
$\textrm{FPR}_{L}(S,\mathcal{D},m)=c(1-e^{-k(S,m)\cdot|S|/m})^{k(S,m)}$
where $c\leq 1$.
We choose realistic values for our example from prior work on evaluating Learned Bloom Filters [KBC+18] on Google’s transparency report. We pick $2$ Megabytes as our memory budget, $m$, chosen from the range of values in Figure 10 of Kraska et al [KBC+18]. We choose the cardinality of the set we want to encode, $|S|$, as $1.7$ million based on the number of unique URLs in Google’s transparency report evaluated in Kraska et al [KBC+18]. With a memory budget of $2$ Megabytes, Kraska et al [KBC+18] demonstrate that a Learned Bloom Filter has $0.25$ of the False Positive Ratio of a Classical Bloom Filter, hence we use that as our value for $c$. We assume $Q_{N}$ to be $0.5$ for this example (results for the complete range of values of $Q_{N}$ can be found in Appendix C).
@@ 5.4 Realistic Example
We use $128$ bits as the size of our security parameter, $\lambda$. For the case of the Downtown Bodega Filter, we let the Bloom Filter Learning Model take $1$ Megabytes, while dividing the remaining $1$ Megabytes equally between Backup Classical Bloom Filters $A$ and $B$. Backup Classical Bloom Filters $A$ and $B$ encode $S^{\prime\prime}$ and $S^{\prime}$ respectively, which are both subsets of the set $S$ (refer to Definition 3.4).
We take $\alpha$ to be a variable ranging from 00 to $1$ equally divided between $\alpha_{P}$ and $\alpha_{N}$ (refer to Appendix C for other strategies of partitioning $\alpha$ between $\alpha_{P}$ and $\alpha_{N}$). Our chosen values are summarized in Table 1. Figure 6 shows the results of our calculations. As can be seen, when the adversary has access to less than a certain cutoff fraction of the workload, the Downtown Bodega Filter outperforms the Secure Classical Bloom Filter for the same memory budget. The $C$ source code for our model and analysis can be found in an anonymously hosted code repository [Bod24]. A thorough experimental evaluation of the Hybrid Model across multiple datasets and a comprehensive range of model parameters can be found in Appendix C.
@@ 6 Open Problems
Problem 1: Secure Constructions for the Learned Bloom Filter in the unsteady setting. While we have provided secure constructions for the Learned Bloom Filter, our results only hold for the steady setting, where the query algorithm $\tilde{Q}_{s}$ may not modify the internal representation of the Learned Bloom Filter. Naor Yogev [NE19] provide secure constructions for the Classical Bloom Filter in an unsteady settings. We leave the formulation of secure constructions for the Learned Bloom Filter in the unsteady setting as an open problem.
Problem 2: Secure Learned Bloom Filters under a stronger adversarial model. The adversarial model we consider for the Learned Bloom Filter is stronger than Naor Yogev [NE19] in that it allows an adversary access to the internal state of the Bloom Filter, but it is weaker than Clayton et al [CPS19] and Filic et al [FPUV22] as it does not allow the adversary to insert entries in the Learned Bloom Filter. We leave the formulation of secure constructions under an adversarial model that allows insertion as an open problem.
Problem 3: Secure Learned Bloom Filters under computationally unbounded adversaries. In this work, we only consider probabilistic polynomial time (PPT) adversaries. An open problem is to prove or disprove any security guarantees for our constructions against computationally unbounded adversaries. Alternatively, an open problem is to provide secure constructions against computationally unbounded adversaries. Naor Yogev [NE19] discuss computationally unbounded adversaries in Section 5 of their work.
7 Acknowledgements
The authors would like to thank Hussain Tirmazi for initial feedback on the paper. We thank the anonymous reviewers for constructive feedback and insightful comments.
@@ 7 Acknowledgements
Appendix A contains discussion of an alternate secure construction of the Learned Bloom Filter which we call the Uptown Bodega Filter. Appendix B contains a brief but self-contained treatment of pseudo-random permutations. Appendix C contains detailed results across a wide range of model parameters for the Hybrid Model. Appendix D contains an overview of our open-source implementation of the Downtown Bodega Filter as well as the Secure Classical Bloom Filter. The appendix also contains results from experiments to investigate how the False Positive Rate of the Downtown Bodega Filter varies compared to the Secure Classical Bloom Filter as we vary the memory budget.
Appendix A The Uptown Bodega Filter
In addition to the Downtown Bodega Filter, discussed in the main text, we consider a second secure construction for the Learned Bloom Filter built on top of the Sandwiched Learned Bloom Filter and a pseudo-random permutation.
Definition A.1 (Uptown).
Let a Uptown Bodega Filter be a data structure $UB_{r,k}=(F_{k},\tilde{C}_{r},\tilde{Q_{s}})$ such that $F_{k}$ is a pseudo-random permutation and $(\tilde{C}_{r},\tilde{Q_{s}})$ is a Sandwiched Learned Bloom Filter representing the set $S^{{}^{\prime}}={\{F_{k}(x):x\in S\}}$ in the steady setting.
Theorem A.2.
Let $SB_{r}$ be an $(n,\epsilon)$-Sandwiched Learned Bloom Filter using $m$ bits of memory. If pseudo-random permutations exist, then there exists a negligible function $\mathrm{negl(\cdot)}$ such that for security parameter $\lambda$ there exists an $(n,\epsilon+\mathrm{negl}(\lambda))$-adversarial resilient Uptown Bodega Filter that uses $m^{{}^{\prime}}=m+\lambda$ bits of memory.
Our proof follows the same line of reasoning as Theorem $4.8$ of Naor Yogev [NE19] which proves that, if pseudo-random permutations exist, any probabilistic polynomial time adversary cannot distinguish between the Classical Bloom Filter with pseudo-random permutations and Challenge 1 by more than a negligible advantage. Due to Lemma 3.2, this result also holds for the Uptown Bodega Filter. The crux of the proof relies on the fact that running $F_{k}$ on an adversary’s queries permits us to consider the queries as random and not chosen adaptively by an adversary, while having no effect on the correctness of the Uptown Bodega Filter.
Proof A.3.
We construct a Uptown Bodega Filter $UB_{k,r}$ from a Sandwiched Bloom Filter $SB_{r}$ as follows. Choose the first key $k\in{\{0,1\}}^{\lambda}$ from pseudo-random permutation, $F_{k}$, over ${\{0,1\}}^{\log|U|}$. Let $S_{k}=F_{k}(S)={\{F_{k}(x):x\in S\}}$. We initialize $SB_{R}$ with $S_{k}$. For each input $x$, our query algorithm outputs $SB_{k,r}(F_{k}(x))$. Just like in the classical Bloom Filter construction of [NE19], the only additional memory required is to store the key $k$ for $F_{k}$, which is $\lambda$ bits. Similarly, the running time of the query algorithm of the Uptown Bodega Filter is one pseudo-random permutation more than the running time of the query algorithm of the Sandwiched Learned Bloom Filter.
The completeness of $UB_{k,r}$ follows from the completeness of $SB_{k,r}$. If $x\in S$ then $SB_{k,r}$ was initialized with $F_{k}(x)$ which will return $1$ from the completeness of $SB_{r}$. The resilience of the construction follows from the following argument: Consider an experiment where $F_{K}$ in the Uptown Bodega Filter is replaced by a truly random permutation oracle $R(\cdot)$. Since $x$ has not been queried, we know that $R(x)$ is a truly random element that was not queried before, and we may think of it as chosen prior to the initialization of $SB_{k,r}$. From the soundness of $SB_{r}$ we get that the probability of $x$ being a false positive is at most $\epsilon$.
Now we show that no probabilistic polynomial time (PPT) adversary $A$ can distinguish between the Uptown Bodega Filter we constructed using $R(\cdot)$ and the Uptown Bodega Filter construction that uses the pseudo-random permutation $F_{k}$ by more than a negligible advantage. Suppose that there does exist a non-negligible function $\delta(\lambda)$ such that $A$ can attack $UB_{k,r}$ and find a false positive with probability $\epsilon+\delta(\lambda)$. We can run $A$ on $UB_{k,r}$ where the oracle is replaced by an oracle that is either random or pseudo-random. We return $1$ if $A$ successfully finds a false positive. This implies that we may distinguish between a truly random permutation and a pseudo-random permutation with probability $\geq\delta(\lambda)$. This contradicts the indistinguishability of pseudo-random permutations.
A.1 Utility of the Uptown Bodega Filter
While we have shown the existence of $(n,\epsilon)$-adversarial resilient Uptown Bodega Filters, it is unclear whether there is any utility in using an Uptown Bodega Filter over a Downtown Bodega Filter or a Secure Classical Bloom Filter. This is due to the fact that we train our learning model on $F_{k}(x)\in S$, a pseudo-random permutation of the set $S$, instead of training it directly on $S$. Intuitively, any structure in $S$ that can be “learned” by a Bloom Filter Learning Model is destroyed by taking the pseudo-random permutation of $S$. In other words, if pseudo-random permutations exist, then to any Bloom Filter Learning Model that runs in Probabilistic Polynomial Time (PPT), $F_{k}(x)\in S$ should appear the same as $|S|$ values sampled from a truly random oracle $R(\cdot)$. A precise proof for this is beyond the scope of this paper.
Appendix B Pseudo-random Permutations
This section provides a brief self-contained treatment of pseudo-random permutations adapted from Chapter 3 of Katz Lindell [KL14] here.

### v4

@@ 4 Discussion
In this section, we discuss the only two known attacks on the Learned Bloom Filter, introduced by Reviriego et al [6]. We refer to Attack 1 from their work as the Blackbox Mutation Attack, and Attack 2 from their work as the Whitebox Mutation Attack respectively. We discuss the Blackbox Mutation Attack in Section 4.1, and the Whitebox Mutation Attack in Section 4.2. Both sections include details on how our Secure Learned Bloom Filter construction mitigates the attack.
@@ 4.1 Black-box Mutation Attack
The black-box adversarial model defined by Reviriego et al [6] is slightly weaker but very similar to the adversarial model we define in Challenge 2.3 of Section 2.2. Both our adversarial model and Reviriego et al’s black-box adversary model allow the adversary access to query the Learned Bloom Filter. One major difference is that our model allows the adversary to choose the initial set $S$ that is represented by the Learned Bloom Filter, whereas Reviriego et al’s model does not. In their attack, Reviriego et al. first test elements until a positive (whether a false positive or true positive) is found. They then mutate the positive by changing a small fraction of the bits in the input to generate more false positives. The attack targets the Bloom Filter Learning Model (recall Figure 2) by making it generate false positives without the input reaching the Backup Bloom Filter. This attack is mitigated by the Downtown Bodega Filter because the Downtown Bodega Filter passes queries through a Secure Classical Backup Bloom Filter even for the case where the Bloom Filter Learning Model returns true i.e when $l(x)>\tau$.
@@ 4.2 White-box Mutation Attack
The white-box adversarial model defined by Reviriego et al [6] is similar to the adversarial model we define in Challenge 2.4 in Section 2.2. With knowledge of the state of the Bloom Filter Learning Model, the adversary can generate mutations in a more sophisticated way. Reviriego et al. provide the example of a malicious URL dataset where an adversary may begin with a non-malicious URL and make changes such as removing the “s” in “https” or removing the “www” to generate false positives. Since we have shown the Downtown Bodega Filter to be $(n,t,\epsilon)$-adversarial reveal resilient in the steady setting, such mutations will not provide the adversary any advantage over the construction.
@@ 5 Hybrid Adversarial Model
In this section, we first define a hybrid model, where part of the queries are chosen by an adversary, while the rest are non-adversarial (“regular”) queries. Next, we analyze the performance of the Downtown Bodega Filter and the Secure Classical Bloom Filter respectively in our hybrid model. We provide results for the conditions in which the Downtown Bodega Filter outperforms the Secure Classical Bloom Filter in the hybrid model. We then discuss the trade-offs of our approach and provide a realistic example where the Downtown Bodega Filter displays better performance as compared to the Secure Classical Bloom Filter under many settings.
@@ 5.1 Hybrid Model
Let $A=(A_{C},A_{Q})$ be a probabilistic polynomial time (PPT) adversary as defined in Section 2. Consider a set of $N$ queries sent to a (Classical or Learned) Bloom Filter. Our adversary, $A$, is allowed to choose exactly $\alpha N$ of those queries, where $\alpha\in[0,1]$. As an example, the $N$ queries may be part of a streaming workload under any of the streaming models described by Muthukrishnan [15].
@@ 5.2 Downtown Bodega Filter
Let $DB_{r}=(\tilde{C}_{r},\tilde{Q}_{s})$ be a Downtown Bodega Filter. Let $l_{r}$ be the corresponding Bloom Filter Learning Model and $\tau$ its threshold. Let $M_{r}^{S^{\prime\prime}}$ be the internal representation of Backup Classical Bloom Filter $A$ and $M_{r}^{S^{\prime}}$ be the internal representation of Backup Classical Bloom Filter $B$. For an adversarial query to generate a false positive in the Downtown Bodega Filter, one of the following must hold for a given query (see Figure 5):
@@
    The query generates a false positive in the Bloom Filter Learning Model and a false positive in the Backup Bloom Filter $A$.
@@
    The query generates a true negative in the Bloom Filter Learning Model and a false positive in the Backup Bloom Filter $B$
@@
Let $\text{FPR}(S,m)$ be the expected false positive probability of a Classical Bloom Filter that encodes the set $S$ with memory budget $m$. Let $\text{FPR}_{L}(S,\mathcal{D},m)$ be the expected false positive probability of a Bloom Filter Learning Model $L$ that encodes the set $S$ using the training dataset $\mathcal{D}$ with memory budget $m$. Similarly, let $\text{TNR}_{L}(S,\mathcal{D},m)$ be the expected true negative probability of a Bloom Filter Learning Model $L$ that encodes the set $S$ using the training dataset $\mathcal{D}$ with memory budget $m$. We assume that the correctness probability of the Bloom Filter Learning Model is independent of the correctness probability of the Backup Classical Bloom Filters. In particular, we assume that for any $m,m^{\prime}$, $\text{FPR}_{L}(m)\cap\text{FPR}(m^{\prime})=\text{FPR}_{L}(m)\text{FPR}(m^{\prime})$ and $\text{TNR}_{L}(m)\cap\text{FPR}(m^{\prime})=\text{TNR}_{L}(m)\text{FPR}(m^{\prime})$.
Consider a system where the total memory budget is $M$. Let the memory allocation of a Downtown Bodega Filter from memory budget $M$ be assigned as follows. Let $m_{L}$ be the number of bits of memory assigned to a Bloom Filter Learning Model $L$. Let $m_{A}$ be the number of bits of memory assigned to Backup Bloom Filter $A$. Let $m_{B}$ be the number of bits of memory assigned to Backup Bloom Filter $B$. Let $\lambda$ be the number of bits assigned to the key of the pseudo-random permutations used before Backup Bloom Filter $A$ and Backup Bloom Filter $B$. Note that to stay within the memory budget it must hold that $M\geq m_{L}+m_{A}+m_{B}+2\lambda$.
Theorem 5.1
@@
Proof
From the definition of a Downtown Bodega Filter (Definition 20, in particular look at the formulation for the query algorithm $\tilde{Q}_{s}$. Figure 5 is also helpful here) it follows that a false positive in the overall construction must either be a false positive in Backup Classical Bloom Filter $A$ or a false positive in Backup Classical Bloom Filter $B$. If the query is a false positive in Backup Classical Bloom Filter $A$, it must also be a false positive in the Bloom Filter Learning Model. Alternatively, if the query is a false positive in Backup Classical Bloom Filter $B$, it must be a true negative in the Bloom Filter Learning Model. The result follows.
@@
Case 1: The query is not adversary-generated. Therefore as established by Theorem 5.1, the False Positive Probability for the query is $\text{FPR}_{DB}(S,\mathcal{D},M)$. There are $(1-\alpha_{P}-\alpha_{N})N$ such queries.
Case 2: The query is adversary-generated such that it generates a false positive in the Bloom Filter Learning Model. Since the Bloom Filter Learning Model generating a false positive and the Bloom Filter Learning Model generating a true negative are mutually exclusive events, the False Positive Probability for the query is the False Positive Probability of Backup Classical Bloom Filter $A$ i.e $\text{FPR}(S^{\prime\prime},m_{A})$. There are $\alpha_{P}N$ such queries.
Case 3: The query is adversary-generated such that it generates a true negative in the Bloom Filter Learning Model. Following logic similar to case 2, we can derive the False Positive Probability of the query to be $\text{FPR}(S^{\prime},m_{B})$. There are $\alpha_{N}N$ such queries.
@@ 5.3 Secure Classical Bloom Filter
An alternative construction is to simply use a well-tuned $(n,t,\epsilon)$-adversarial reveal resilient Classical Bloom Filter. We will refer to this construction in this section as the Secure Classical Bloom Filter without confusion. The expected false positive probability of the Secure Classical Bloom Filter encoding a set $S$ is, by definition, $\text{FPR}(S,m_{L}+m_{A}+m_{B}+\lambda)$. The extra $\lambda$ is because the Secure Classical Bloom Filter requires one less pseudo-random permutation than the Downtown Bodega Filter.
@@
Theorem 5.3
For a given memory budget $M=m_{L}+m_{A}+m_{B}+2\lambda$, any set $S$ and corresponding training dataset $\mathcal{D}$, in the hybrid adversarial setting with given $\alpha=\alpha_{P}+\alpha_{N}$, the expected false positive probability of the Downtown Bodega Filter is lower than the expected false positive probability of the Secure Classical Bloom Filter if the following holds:
@@
5.4 Realistic Example
Mitzenmacher and Broder [4] show that the false positive probability for a Classical Bloom Filter with $m$ bits encoding a set $S$, using $k(S,m)$ hash functions is
$\textrm{FPR}(S,m)=(1-e^{-k(S,m)\cdot|S|/m})^{k(S,m)}$
In our analysis, we use the value of the number of hash functions $k(S,m)$ is always optimally chosen to be $k(S,m)=\ln 2\cdot(m/|S|)$ (this optimal value is also derived by Mitzenmacher and Broder [4]). We further model the false positive rate of a Bloom Filter Learning Model as being the same as the false positive rate of a Classical Bloom Filter encoding set $S$ but with a better false positive probability for the same memory budget. This is consistent with the assumptions made by prior work including Kraska et al [5] and Mitzenmacher [7].
$\textrm{FPR}_{L}(S,\mathcal{D},m)=c(1-e^{-k(S,m)\cdot|S|/m})^{k(S,m)}$
where $c\leq 1$.
We choose realistic values for our example from prior work on evaluating Learned Bloom Filters [5] on Google’s transparency report. We pick $2$ Megabytes as our memory budget, $m$, chosen from the range of values in Figure 10 of Kraska et al [5]. We choose the cardinality of the set we want to encode, $|S|$, as $1.7$ million based on the number of unique URLs in Google’s transparency report evaluated in Kraska et al [5]. With a memory budget of $2$ Megabytes, Kraska et al [5] demonstrate that a Learned Bloom Filter has $0.25$ of the False Positive Ratio of a Classical Bloom Filter, hence we use that as our value for $c$. We assume $Q_{N}$ to be $0.5$ for this example (results for the complete range of values of $Q_{N}$ can be found in Appendix 0.B).
We use $128$ bits as the size of our security parameter, $\lambda$. For the case of the Downtown Bodega Filter, we let the Bloom Filter Learning Model take $1$ Megabytes while dividing the remaining $1$ Megabytes equally between Backup Classical Bloom Filters $A$ and $B$. Backup Classical Bloom Filters $A$ and $B$ encode $S^{\prime\prime}$ and $S^{\prime}$ respectively, which are both subsets of the set $S$ (refer to Definition 20).
We take $\alpha$ to be a variable ranging from 00 to $1$ equally divided between $\alpha_{P}$ and $\alpha_{N}$ (refer to Appendix 0.B for other strategies of partitioning $\alpha$ between $\alpha_{P}$ and $\alpha_{N}$). Our chosen values are summarized in Table 1. Figure 6 shows the results of our calculations. As can be seen, when the adversary has access to less than a certain cutoff fraction of the workload, the Downtown Bodega Filter outperforms the Secure Classical Bloom Filter for the same memory budget. The $C$ source code for our model and analysis can be found in an anonymously hosted code repository [16]. A thorough experimental evaluation of the Hybrid Model across multiple datasets and a comprehensive range of model parameters can be found in Appendix 0.B.
@@ 6 Open Problems
Problem 1: Security in the unsteady setting. While we have provided secure constructions for the Learned Bloom Filter, our results only hold for the steady setting, where the query algorithm $\tilde{Q}_{s}$ may not modify the internal representation of the Learned Bloom Filter. Naor and Yogev [8] provide secure constructions for the Classical Bloom Filter in an unsteady setting. We leave the formulation of secure constructions for the Learned Bloom Filter in the unsteady setting as an open problem.
Problem 2: Security in a stronger adversarial model. The adversarial model we consider for the Learned Bloom Filter is stronger than Naor and Yogev [8] in that it allows an adversary access to the internal state of the Bloom Filter, but it is weaker than the Bet-Or-Pass (BP) test security notion Naor and Oved [11] as proven in their work. We leave the construction of a Bet-Or-Pass test resilient Learned Bloom Filter as an open problem.
Problem 3: Security when allowing insertions Clayton et al. [12] and Filic et al. [13] prove security results for an adversary that can not only query but also insert entries in the Learned Bloom Filter. We leave the formulation of secure constructions of the Learned Bloom Filter under an adversarial model that allows insertion as an open problem.
Problem 3: Security against computationally unbounded adversaries. In this work, we only consider probabilistic polynomial time (PPT) adversaries. An open problem is to prove or disprove any security guarantees for our constructions against computationally unbounded adversaries. Alternatively, an open problem is to provide secure constructions against computationally unbounded adversaries. Naor and Yogev [8] discuss computationally unbounded adversaries in Section 5 of their work.
{credits}
[…]
The authors have no competing interests
@@
Appendix 0.A contains a brief but self-contained treatment of pseudo-random permutations. Appendix 0.B contains detailed results across a wide range of model parameters for the Hybrid Model.
Appendix 0.A Pseudo-random Permutations
This section provides a brief self-contained treatment of pseudo-random permutations adapted from Chapter 3 of Katz Lindell [14] here.